16 lbs of Brand M Cinnamon was made by combining 12 lbs of Indonesian cinnamon which costs $19/ lb with 4 lb of Thai cinnamon which costs $11/ lb. Find the cost per lb of the mixture.
Answer by TimothyLamb(4379)

i = weight of Indonesian cinnamon = 12 lbs
t = weight of Thai cinnamon = 4 lbs
---
19i + 11t = 16*C
C = cost per pound of mixture
---
19i + 11t = 16*C
C = (19i + 11t)/16
C = (19*12 + 11*4)/16
C = (228 + 44)/16
---
C = 17 dollars/lb
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php